OpenStudy (anonymous):

find the y intercept and equation of the line passing through points P (-2,6) and R (5,0)

OpenStudy (anonymous):

equation for a line is\[y=mx+b\]Plugging in each point you get\[6=2m+b\]and\[0=5m+b\]So\[b=6-2m=-5m\]\[m=-2\]\[b=12\]b is the y intercept and teh equation is \[y=-2x+12\]

OpenStudy (asadkarim7):

gradient = 0-6/5-(-2) = -6 / 7 y -6= -6/7 ( x-(-2)) 7y -42 = -6x -12 7y = -6x +30

OpenStudy (anonymous):

why did you put -6 next to the y and how did you come up with (x-(-2)?? im use to the equation y=mx+b...... which is why im a bit confused... (you did get the answer right though).

OpenStudy (asadkarim7):

oh lets do it your way y= mx+c 6 = -6/7 (-2) + c 6 -12/7 =c -5 / 7 = c y = -6/7 x -5/7 7y = -6x - 5

OpenStudy (anonymous):

but then the answer comes out wrong... the -5 has to be a +30 hahahaha... im sorry im just really bad at math... the answer is y=-6/7x+30/7 or 7y+6x=30 which is what you had the first time.

OpenStudy (asadkarim7):

oh maybe i'have made mistake ist time

OpenStudy (asadkarim7):

try yourself and tell me what answer are u getting

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

omg thank you soo much, I figured out what i was doing wrong thanks to your example... Thank You :)

OpenStudy (asadkarim7):

and what was i doing wrong ex 1 or example 2

OpenStudy (anonymous):

example 2... but just one little mistake. instead of -5/7 it's 30/7. b=30/7 making the equation y=-6/7x+30/7 :D

OpenStudy (asadkarim7):

ohk btw im watching movie so lost concentration :P

OpenStudy (anonymous):

hahahaha its ok... what movie?

OpenStudy (asadkarim7):

pirates of carribean 4

OpenStudy (anonymous):

can you help me with another question by the way??? it has to do with the same points and numbers... :/

OpenStudy (asadkarim7):

oh k

OpenStudy (asadkarim7):

post it

OpenStudy (anonymous):

how do you find the slope and y intercept for the equation 2y+x=6?

OpenStudy (asadkarim7):

ok the slope can be formed by arranging the equation in the form y = mx+c in this case 2y = -x +6 y = -1/2 x + 6/2 y = -1/2x +3 the slope is -1/2 and y intercept is 3

OpenStudy (anonymous):

thank you :D your smart

OpenStudy (asadkarim7):

welcome dude :D